OpenStudy (anonymous):

Samantha puts an initial $500 into a savings account. The account has a 4% annual compound interest rate. What is the function that best represents Samantha's account and how much does she have in savings after 5 years? Round your answer to the nearest dollar. Hint: Use the formula, f(x) = P(1 + r)x.

OpenStudy (anonymous):

f(x) = 500(1.4)x, $2,689 f(x) = 500(0.04)x, $512 f(x) = 500(1.04)x, $608 f(x) = 500(1.04)x, $5

OpenStudy (anonymous):

I will fan and give a medal please help me

OpenStudy (anonymous):

f(x) = 500(1.4)x, $2,689 f(x) = 500(0.04)x, $512 f(x) = 500(1.04)x, $608 f(x) = 500(1.04)x, $5

OpenStudy (anonymous):

Are these given answers to choose from?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But I have no idea where to start

OpenStudy (anonymous):

First I thought that it would be f(x)=500(40%)^5

OpenStudy (anonymous):

Okay well, lets think of what we already know. Using the hint and the given answers we already know that P represents the amount Samantha put in already. Which is $500 right?

OpenStudy (anonymous):

Now in the hint it shows to add R to 1. What do you think R represents?

OpenStudy (anonymous):

Are you sure that in the parenthesis of hint you add 1 to r?

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

What is r

OpenStudy (mathstudent55):

Start with P amount and r interest rate. r is written as a decimal. For example, if the interest rate is 5%, you write it as 0.05 After 1 year, you have P(1 + r) After 2 years , you have P(1 + r)(1 + r) = P(1 + r)^2 After 3 years, you have P(1 + r)^2(1 + r) = P(1 + r)^3 After 4 years, you have P(1 + r)^3(1 + r) = P(1 + r)^4 If you follow this pattern, after x years, you have P(1 + r)^x Notice that this is the formula the problem gives you in the hint. Now replace P with 500 and r with 0.04, but leave x in. What formula do you have now? Finally use the new formula with x = 5 and evaluate it.

OpenStudy (anonymous):

Thank you so much foryour help #mathstudent55 how do i give you a medal

OpenStudy (mathstudent55):

You're welcome. To give a medal, click on the blue box "Best Response" of the response you want to give a medal to.